# Diagonal Simplicial Tensor Modules and Algebraic $n$-Hypergroupoids

### v1

@@ Abstract
Let $A$ be a commutative ring, $k\in\mathbb{Z}^{+}$, and $\vec{s}=(n_{1},\dots,n_{k})\in(\mathbb{Z}^{+})^{k}$ with $n=\min(\vec{s})-1$. We attach to $\vec{s}$ a diagonal simplicial tensor module $X_{\bullet}(\vec{s};A)$ whose $p$-simplices are functions on a cosimplicial index set $I_{p}(\vec{s})\subseteq\mathbb{N}^{k}$. This extends Quillen’s diagonal on double simplicial groups to $k$ commuting directions in $\mathbf{Mod}_{A}$, compatibly with Dold–Kan normalization.
We analyze the horn kernels $R_{p,j}(X)$ via “missing indices” and show that $R_{p,j}(X)\neq 0$ if and only if $k\geq p$, independently of $j$. It follows that $X_{\bullet}(\vec{s};A)$ is an algebraic $n$-hypergroupoid in the sense of Duskin–Glenn precisely when $k\leq n$, and that horn fillers in dimension $n$ are non-unique precisely when $k\geq n$; in particular $X_{\bullet}(\vec{s};A)$ is strict in our sense precisely when $k=n$. A Horn Non-Degeneracy Lemma then shows that $R_{n,j}(X)\cap D_{n}=\{0\}$ and yields a decomposition $X_{n}=R_{n,j}(X)\oplus D_{n}$. A shift-and-truncate chain homotopy, $\operatorname{Stab}(\vec{s})$-equivariant and filtration-preserving, shows that $X_{\bullet}(\vec{s};A)$ is contractible and forces the associated spectral sequence to collapse at $E_{1}$.
When $A$ is an infinite field $K$, we classify simplicial submodules generated by a single tensor via kernel sequences and a moduli map to a product of Grassmannians, obtaining an irreducible and unirational incidence variety.
@@ 1 Introduction
This construction generalizes Quillen’s diagonal on a double simplicial group [Quillen1966] to $k$ commuting simplicial directions in the category $\mathbf{Mod}_{A}$ and is compatible with Dold–Kan normalization in the sense of Dold, Kan, May, Weibel, and Goerss–Jardine [Dold1958, Kan1958, MaySimplicial, Weibel, GoerssJardine].
@@
and show that their combinatorics controls the homotopy-theoretic behavior of the DSTM. For the simplicial dimension $n=\min(\vec{s})-1$ we prove that $X_{\bullet}(\vec{s};A)$ is a strict algebraic $n$-hypergroupoid in this sense if and only if the tensor order satisfies $k=n$. The transition from non-unique to unique horn fillers occurs exactly at $p=n$, and this threshold is determined by the tensor order.
@@
In the abelian setting of simplicial modules, horn fillers always exist (the Kan condition is automatic). The obstruction to the uniqueness of fillers is captured by the horn kernel $R_{p,j}$. Following terminology from Duskin–Glenn [Duskin1979, Glenn1982] and the nnLab, a simplicial module is a strict algebraic $n$-hypergroupoid if fillers are unique in dimensions strictly greater than $n$ ($R_{p,j}=0$ for $p>n$), and non-unique in dimension $n$ ($R_{n,j}\neq 0$).
@@
$K(T)_{\bullet}:=(\ker f_{T,p})_{p}\subseteq K[\Delta^{n}]_{\bullet},$
@@
We define a moduli map $\Psi$ from a Zariski-open locus $\mathcal{U}\subset X_{n}(\vec{s};K)$ to a product of Grassmannians by sending $T$ to the collection of subspaces $K(T)_{p}\subseteq K[\Delta^{n}]_{p}$. The image $\mathcal{M}(\vec{s})$ lies within a closed incidence subvariety defined by simplicial compatibility conditions, which we show are linear in the Segre–Plücker coordinates. This reduces the classification of these generated submodules to the geometry of these incidence conditions and of the index collision maps $\mathcal{I}_{p}$.
@@
Section 2 establishes notation and the DSTM construction. Section 3 characterizes missing indices and the horn kernel basis. Section 4 develops the Moore filler algorithm and normalization. Section 5 establishes rank formulas and the classification theorem. Section 6 proves the Horn Non-Degeneracy Lemma and the decomposition $X_{n}=R_{n,j}\oplus D_{n}$. Section 7 develops the homology dichotomy and the hypergroupoid classification. Section 8 develops the algebraic geometry of generated submodules. A proves contractibility via an explicit equivariant homotopy.
@@ 2.1 Simplicial preliminaries
The simplicial category $\Delta$ has objects $p]:=\{0,\ldots,p\}$ for $p\in\mathbb{N}$; morphisms are nondecreasing maps. It is generated by the coface maps $\delta_{i}^{p}:[p-1]\to[p]$ (the injection missing $i$) and the codegeneracy maps $\sigma_{i}^{p}:[p+1]\to[p]$ (the surjection repeating $i$).
A simplicial $A$-module is a functor $X_{\bullet}:\Delta^{\mathrm{op}}\to\mathbf{Mod}_{A}$. We write $X_{p}:=X([p])$. The induced maps are the face maps $d_{i}^{p}:=X(\delta_{i}^{p}):X_{p}\to X_{p-1}$ and the degeneracy maps $s_{i}^{p}:=X(\sigma_{i}^{p}):X_{p}\to X_{p+1}$. The associated chain complex $(X_{\bullet},\partial_{\bullet})$ has boundary maps $\partial_{p}=\sum_{i=0}^{p}(-1)^{i}d_{i}^{p}$.
@@ 2.3 The Cosimplicial Index Set $I_{\bullet}$
We define a functor $I_{\bullet}:\Delta\to\mathbf{Set}$. On objects, $I_{\bullet}([p]):=I_{p}$. Note that $M_{a}(p\pm 1)=M_{a}(p)\pm 1$.
The structure maps are defined ”diagonally” as products of the standard generators.
@@
The functor $I_{\bullet}$ maps the generators of $\Delta$ as follows:
@@ 2.4 The Diagonal Simplicial Module $X_{\bullet}(\vec{s};A)$
The diagonal simplicial $A$-tensor module $X_{\bullet}(\vec{s};A):\Delta^{\mathrm{op}}\to\mathbf{Mod}_{A}$ is the composition $A^{(-)}\circ I_{\bullet}$.
@@
so an $N\times N$ matrix may be viewed as a simplex in degree N1N-1 of $X_{\bullet}(\vec{s};A)$. This identification will be used when we regard adjacency matrices of graphs on $N$ vertices as elements of $X_{N-1}(\vec{s};A)$.
@@ 2.6 Base Change and Realization Maps
Let $T\in X_{n}(\vec{s};A)$, and let $f_{T}:A[\Delta^{n}]\to X_{\bullet}(\vec{s};A)$ be the corresponding realization map (defined in Section 8). Let $\varphi:A\to B$ be a ring homomorphism, and let $T_{B}:=\theta_{n}(T\otimes 1)\in X_{n}(\vec{s};B)$. Let $\varphi_{p}:A[\Delta^{n}]_{p}\otimes_{A}B\xrightarrow{\cong}B[\Delta^{n}]_{p}$ denote the canonical isomorphism.
@@
    $A[\Delta^{n}]_{p}\otimes_{A}B$$X_{p}(\vec{s};A)\otimes_{A}B$$B[\Delta^{n}]_{p}$$X_{p}(\vec{s};B)$$\scriptstyle{f_{T,p}\otimes 1_{B}}$$\scriptstyle{\varphi_{p}}\scriptstyle{\cong}$$\scriptstyle{\theta_{p}}\scriptstyle{\cong}$$\scriptstyle{f_{T_{B},p}}$
@@
(1) Let $\alpha\in\Delta([p],[n])$ be a basis element of $A[\Delta^{n}]_{p}$. Both paths in the diagram map the element $\alpha\otimes b$ to $b\cdot f_{T_{B},p}(\alpha)$, establishing commutativity.
(2) Consider the exact sequence $0\to\ker f_{T,p}\to A[\Delta^{n}]_{p}\xrightarrow{f_{T,p}}X_{p}(\vec{s};A)$. If $B$ is a flat $A$-module, the functor $(-)\otimes_{A}B$ is exact. Applying it yields the exact sequence
$0\to(\ker f_{T,p})\otimes_{A}B\to A[\Delta^{n}]_{p}\otimes_{A}B\xrightarrow{f_{T,p}\otimes 1_{B}}X_{p}(\vec{s};A)\otimes_{A}B.$
@@ 3.2 Structure of the Face Kernels
since $d_{i}^{p}(E_{m})=0$ if $m\notin\operatorname{im}(\Delta_{i}^{p})$ (Lemma 3.3). By the Injectivity Lemma 3.4, the set $\{d_{i}^{p}(E_{m})\mid m\in\operatorname{im}(\Delta_{i}^{p})\}$ consists of distinct standard basis elements in $X_{p-1}$, hence is AA-linearly independent.
Thus $d_{i}^{p}(T)=0$ if and only if $a_{m}=0$ for all $m\in\operatorname{im}(\Delta_{i}^{p})$. ∎
@@ 3.3 Missing Indices and the Support Characterization
By Theorem 3.5, $\ker(d_{i}^{p})$ is the coordinate subspace spanned by the basis $B_{i}=\{E_{m}\mid m\notin\operatorname{im}(\Delta_{i}^{p})\}$. The horn kernel $R_{p,j}$ is the intersection of these coordinate subspaces for $i\in F_{j}$. The intersection of coordinate subspaces is spanned by the intersection of their bases, $\bigcap_{i\in F_{j}}B_{i}$. An element $E_{m}$ lies in this intersection if and only if $m\notin\operatorname{im}(\Delta_{i}^{p})$ for all $i\in F_{j}$, which defines $M_{p,j}$. The consequence regarding the support of the difference follows because $T_{1}-T_{2}\in R_{p,j}$. ∎
@@ 3.4 Combinatorial Characterization
($\Leftarrow$) Suppose $k\geq p$. Let $F_{j}=\{f_{1},\dots,f_{p}\}$. Define $m=(f_{1},\dots,f_{p},0,\dots,0)\in\mathbb{N}^{k}$. Since $M_{a}(p)\geq p$, we have $p]^{k}\subseteq I_{p}$, so $m\in I_{p}$. Since $\operatorname{im}(m)\supseteq F_{j}$, $m\in M_{p,j}$. ∎
@@ 3.6 Genericity
Let $T^{\prime}=\mu_{j}(\Phi_{j}(T))$ be the Moore filler of $T$ (see Section 4). The difference $T-T^{\prime}$ lies in $R_{p,j}$. By Theorem 3.7, $\operatorname{supp}(T-T^{\prime})\subseteq M_{p,j}$.
@@
Assume $A$ is an integral domain. A tensor $T\in X_{p}$ is called $(p,j)$-generic if the difference $T-T^{\prime}$ has maximal support: $\operatorname{supp}(T-T^{\prime})=M_{p,j}$. (This requires $k\geq p$).
@@
Assume $A$ is an integral domain and $k\geq p$. The set of $(p,j)$-generic tensors forms a Zariski-open subset $\mathcal{U}_{p,j}\subset X_{p}$. If moreover $A$ is infinite, then $\mathcal{U}_{p,j}$ is non-empty.
@@
Let $R(T)=T-T^{\prime}$. The map $T\mapsto T^{\prime}$ is $A$-linear (as $\mu_{j}$ and $\Phi_{j}$ are linear), hence $R(T)$ is $A$-linear. Write $R(T)=\sum_{m\in M_{p,j}}c_{m}(T)E_{m}$. Each coefficient $c_{m}(T)$ is a linear form in the entries of $T$.
@@
The complement is the union of the zero loci of the non-zero forms $c_{m}$, hence a finite union of proper linear subspaces. Therefore $\mathcal{U}_{p,j}$ is Zariski-open. If $A$ is infinite, a finite union of proper linear subspaces cannot equal $X_{p}$, so $\mathcal{U}_{p,j}$ is non-empty. ∎

### v2

@@ Abstract
Let $A$ be a commutative ring, let $k\in\mathbb{Z}^{+}$, and let $\vec{s}=(n_{1},\dots,n_{k})\in(\mathbb{Z}^{+})^{k}$ with $n=\min_{a}(n_{a})-1$. We attach to $\vec{s}$ a diagonal simplicial tensor module $X_{\bullet}(\vec{s};A)$ whose $p$-simplices are functions on a cosimplicial index set $I_{p}(\vec{s})\subseteq\mathbb{N}^{k}$. This extends Quillen’s diagonal on double semi-simplicial groups: $X_{\bullet}(\vec{s};A)$ is obtained by restricting a $k$-fold simplicial $A$-module along the diagonal $p\mapsto(p,\ldots,p)$.
Using a “missing indices” description of face kernels, we compute the horn kernels $R_{p,j}(X)$ and show that $R_{p,j}(X)\neq 0$ if and only if $k\geq p$, independently of $j$. Consequently, $X_{\bullet}(\vec{s};A)$ is an algebraic $n$-hypergroupoid in the sense of Duskin (1979) and Glenn (1982) if and only if $k\leq n$, and horn fillers in dimension $n$ are non-unique if and only if $k\geq n$; in particular it is strict precisely when $k=n$. A Horn Non-Degeneracy Lemma shows that, for $p\geq 1$, $R_{p,j}(X)\cap D_{p}(X)=\{0\}$ and yields a decomposition $X_{p}=R_{p,j}(X)\oplus D_{p}(X)$. An explicit shift-and-truncate chain homotopy, equivariant under $\operatorname{Stab}(\vec{s})$ and compatible with a natural filtration, contracts $X_{\bullet}(\vec{s};A)$ and forces the associated spectral sequence to collapse at $E_{1}$.
When $A$ is an infinite field $K$, we study simplicial submodules generated by a single tensor via kernel sequences and a moduli map to a product of Grassmannians. The moduli map image is an irreducible and unirational constructible subset of a determinantal incidence variety.
@@ 1 Introduction
This construction generalizes Quillen’s diagonal on a double semi-simplicial group [6]: it is the restriction $p\mapsto(p,\ldots,p)$ of a $k$-fold simplicial $A$-module in $\mathbf{Mod}_{A}$ (Remark 2.3), and is compatible with Dold–Kan normalization [1, 4, 8].
@@
and show that their combinatorics controls the homotopy-theoretic behavior of the DSTM. For the simplicial dimension $n=\min(\vec{s})-1$ we prove that $X_{\bullet}(\vec{s};A)$ is a strict algebraic $n$-hypergroupoid in this sense if and only if the tensor order satisfies $k=n$. The transition from non-unique to unique horn fillers occurs at $p=k$ (hence at $p=n$ in the strict case $k=n$).
@@
In the abelian setting of simplicial modules, horn fillers always exist (the Kan condition is automatic). The obstruction to the uniqueness of fillers is captured by the horn kernel $R_{p,j}$. Following terminology from Duskin and Glenn [2, 3], a simplicial module is a strict algebraic $n$-hypergroupoid if fillers are unique in dimensions strictly greater than $n$ ($R_{p,j}=0$ for $p>n$), and non-unique in dimension $n$ ($R_{n,j}\neq 0$).
@@
$K(T)_{\bullet}:=(\ker f_{T,p})_{p}\subseteq K[\boldsymbol{\Delta}^{n}]_{\bullet},$
@@
We define a moduli map $\Psi$ from a Zariski-open locus $\mathcal{U}\subset X_{n}(\vec{s};K)$ to a product of Grassmannians by sending $T$ to the collection of subspaces $K(T)_{p}\subseteq K[\boldsymbol{\Delta}^{n}]_{p}$. The image $\mathcal{M}(\vec{s})$ lies within a closed incidence subvariety defined by simplicial compatibility conditions, which we show are linear in the Segre–Plücker coordinates. This reduces the classification of these generated submodules to the geometry of these incidence conditions and of the index collision maps $\mathcal{I}_{p}$.
@@
Section 2 establishes notation and the DSTM construction. Section 3 characterizes missing indices and the horn kernel basis. Section 4 develops the Moore filler algorithm and normalization. Section 5 establishes rank formulas and the classification theorem. Section 6 proves the Horn Non-Degeneracy Lemma and the decomposition $X_{n}=R_{n,j}\oplus D_{n}$. Section 7 develops the homology dichotomy and the hypergroupoid classification. Section 8 develops the algebraic geometry of generated submodules. In A we prove contractibility via an explicit equivariant homotopy.
@@ 2.1 Simplicial preliminaries
The simplicial category $\boldsymbol{\Delta}$ has objects $p]:=\{0,\ldots,p\}$ for $p\in\mathbb{N}$; morphisms are nondecreasing maps. It is generated by the coface maps $\delta_{i}^{p}:[p-1]\to[p]$ (the injection missing $i$) and the codegeneracy maps $\sigma_{i}^{p}:[p+1]\to[p]$ (the surjection repeating $i$).
A simplicial $A$-module is a functor $X_{\bullet}:\boldsymbol{\Delta}^{\mathrm{op}}\to\mathbf{Mod}_{A}$. We write $X_{p}:=X([p])$. The induced maps are the face maps $d_{i}^{p}:=X(\delta_{i}^{p}):X_{p}\to X_{p-1}$ and the degeneracy maps $s_{i}^{p}:=X(\sigma_{i}^{p}):X_{p}\to X_{p+1}$. The associated chain complex $(X_{\bullet},\partial_{\bullet})$ has boundary maps $\partial_{p}=\sum_{i=0}^{p}(-1)^{i}d_{i}^{p}$.
@@ 2.3 The Cosimplicial Index Set $I_{\bullet}$
We define a functor $I_{\bullet}:\boldsymbol{\Delta}\to\mathbf{Set}$. On objects, $I_{\bullet}([p]):=I_{p}$. Note that $M_{a}(p\pm 1)=M_{a}(p)\pm 1$.
The structure maps are defined by applying the same simplicial generator in each tensor axis.
@@
The functor $I_{\bullet}$ maps the generators of $\boldsymbol{\Delta}$ as follows:
@@ 2.4 The Diagonal Simplicial Module $X_{\bullet}(\vec{s};A)$
The diagonal simplicial $A$-tensor module $X_{\bullet}(\vec{s};A):\boldsymbol{\Delta}^{\mathrm{op}}\to\mathbf{Mod}_{A}$ is the composition $A^{(-)}\circ I_{\bullet}$.
@@
Remark 2.3.
Define $\widetilde{X}:(\boldsymbol{\Delta}^{\mathrm{op}})^{k}\to\mathbf{Mod}_{A}$ by
$\widetilde{X}_{p_{1},\ldots,p_{k}}:=A^{\prod_{a=1}^{k}[\,M_{a}(p_{a})\,]}.$
In the $a$-th simplicial direction, the face and degeneracy maps are induced by precomposition with $\delta_{i}^{M_{a}(p_{a})}$ and $\sigma_{i}^{M_{a}(p_{a})}$ on the $a$-th factor (identity on the others). Then
$X_{p}(\vec{s};A)=\widetilde{X}_{p,\ldots,p},$
so the structure maps of $X_{\bullet}(\vec{s};A)$ are obtained by applying the same simplicial operator in each direction.
@@
so an $N\times N$ matrix may be viewed as a simplex in degree N1N-1 of $X_{\bullet}(\vec{s};A)$.
@@ 2.6 Base Change and Realization Maps
Let $T\in X_{n}(\vec{s};A)$, and let $f_{T}:A[\boldsymbol{\Delta}^{n}]\to X_{\bullet}(\vec{s};A)$ be the corresponding realization map (defined in Section 8). Let $\varphi:A\to B$ be a ring homomorphism, and let $T_{B}:=\theta_{n}(T\otimes 1)\in X_{n}(\vec{s};B)$. Let $\varphi_{p}:A[\boldsymbol{\Delta}^{n}]_{p}\otimes_{A}B\xrightarrow{\cong}B[\boldsymbol{\Delta}^{n}]_{p}$ denote the canonical isomorphism.
@@
    $A[\boldsymbol{\Delta}^{n}]_{p}\otimes_{A}B$$X_{p}(\vec{s};A)\otimes_{A}B$$B[\boldsymbol{\Delta}^{n}]_{p}$$X_{p}(\vec{s};B)$$\scriptstyle{f_{T,p}\otimes 1_{B}}$$\scriptstyle{\varphi_{p}}\scriptstyle{\cong}$$\scriptstyle{\theta_{p}}\scriptstyle{\cong}$$\scriptstyle{f_{T_{B},p}}$
@@
(1) Let $\alpha\in\boldsymbol{\Delta}([p],[n])$ be a basis element of $A[\boldsymbol{\Delta}^{n}]_{p}$. Both paths in the diagram map the element $\alpha\otimes b$ to $b\cdot f_{T_{B},p}(\alpha)$, establishing commutativity.
(2) Consider the exact sequence $0\to\ker f_{T,p}\to A[\boldsymbol{\Delta}^{n}]_{p}\xrightarrow{f_{T,p}}X_{p}(\vec{s};A)$. If $B$ is a flat $A$-module, the functor $(-)\otimes_{A}B$ is exact. Applying it yields the exact sequence
$0\to(\ker f_{T,p})\otimes_{A}B\to A[\boldsymbol{\Delta}^{n}]_{p}\otimes_{A}B\xrightarrow{f_{T,p}\otimes 1_{B}}X_{p}(\vec{s};A)\otimes_{A}B.$
@@ 3.2 Structure of the Face Kernels
since $d_{i}^{p}(E_{m})=0$ if $m\notin\operatorname{im}(\Delta_{i}^{p})$ (Lemma 3.3). By the Injectivity Lemma 3.4, the set $\{d_{i}^{p}(E_{m})\mid m\in\operatorname{im}(\Delta_{i}^{p})\}$ consists of distinct standard basis elements in $X_{p-1}$, hence is $A$-linearly independent.
Thus $d_{i}^{p}(T)=0$ if and only if $a_{m}=0$ for all $m\in\operatorname{im}(\Delta_{i}^{p})$. Equivalently, $T$ is an $A$-linear combination of the $E_{m}$ with $m\notin\operatorname{im}(\Delta_{i}^{p})$, which proves the stated basis description of $\ker(d_{i}^{p})$. ∎
@@ 3.3 Missing Indices and the Support Characterization
By Theorem 3.5, $\ker(d_{i}^{p})$ is the coordinate subspace spanned by the basis $B_{i}=\{E_{m}\mid m\notin\operatorname{im}(\Delta_{i}^{p})\}$. The horn kernel $R_{p,j}$ is the intersection of these coordinate subspaces for $i\in F_{j}$. Since each $\operatorname{span}_{A}(B_{i})$ consists exactly of those tensors whose coefficients vanish outside $B_{i}$, their intersection consists of those tensors whose coefficients vanish outside $\bigcap_{i\in F_{j}}B_{i}$, i.e. it is $\operatorname{span}_{A}\!\bigl(\bigcap_{i\in F_{j}}B_{i}\bigr)$.
An element $E_{m}$ lies in this intersection if and only if $m\notin\operatorname{im}(\Delta_{i}^{p})$ for all $i\in F_{j}$, which defines $M_{p,j}$. The consequence regarding the support of the difference follows because $T_{1}-T_{2}\in R_{p,j}$. ∎
@@ 3.4 Combinatorial Characterization
($\Leftarrow$) Suppose $k\geq p$. Let $F_{j}=\{f_{1},\dots,f_{p}\}$. Set
$m:=(f_{1},\dots,f_{p},0,\dots,0)\in[p]^{k}\subseteq I_{p}.$
Then $\operatorname{im}(m)\supseteq F_{j}$, so $m\in M_{p,j}$. ∎
@@ 3.6 Genericity
Fix $p\geq 1$ and $0\leq j\leq p$, and assume $A$ is an infinite field and $k\geq p$.
For $T\in X_{p}$, set
$T^{\prime}:=\mu_{j}(\Phi_{j}(T)),$
the Moore filler of the horn $\Phi_{j}(T)$ (Section 4). Then $\Phi_{j}(T^{\prime})=\Phi_{j}(T)$, hence $T-T^{\prime}\in R_{p,j}$. By Theorem 3.7, $\operatorname{supp}(T-T^{\prime})\subseteq M_{p,j}$.
@@
A tensor $T\in X_{p}$ is called $(p,j)$-generic if the difference $T-T^{\prime}$ has maximal support: $\operatorname{supp}(T-T^{\prime})=M_{p,j}$. (This requires $k\geq p$).
@@
Assume $A$ is an infinite field and $k\geq p$. The set of $(p,j)$-generic tensors forms a Zariski-open subset $\mathcal{U}_{p,j}\subset X_{p}$.
@@
Define the $A$-linear endomorphism
$R:=\operatorname{id}_{X_{p}}-\mu_{j}\circ\Phi_{j}:X_{p}\to X_{p}.$
Then $R(T)=T-\mu_{j}(\Phi_{j}(T))=T-T^{\prime}\in R_{p,j}$. Write
$R(T)=\sum_{m\in M_{p,j}}c_{m}(T)\,E_{m}.$
Each coefficient $c_{m}(T)$ is an $A$-linear form in the entries of $T$.
@@
The complement is the union of the zero loci of the non-zero forms $c_{m}$, hence a finite union of proper linear subspaces. Therefore $\mathcal{U}_{p,j}$ is Zariski-open. ∎